If 64^2n+1 = 16^4n-1, what is the value of n?
Answer by Alan3354(69443)

If 64^2n+1 = 16^4n-1, what is the value of n?
---
If you mean If 64^(2n+1) = 16^(4n-1), what is the value of n?
(4^3)^(2n+1) = (4^2)^(4n-1)
4^(6n+3) = 4^(8n-2)
6n+3 = 8n-2
n = 5/2